The half- life of a radioactive substance is one hundred ninety-four days. How many days will it take for eighty percent of the substance to decay?
I used the half life decay formula.
my formula for the problem read...
0.8Nsubscript0=Nsubscript0*(1/2)superscript(t/194)
the result was about 62.5,
is this correct.
=]

Answer by ankor@dixie-net.com(12706)

The half- life of a radioactive substance is one hundred ninety-four days. How many days will it take for eighty percent of the substance to decay?
I used the half life decay formula.
my formula for the problem read...
0.8Nsubscript0=Nsubscript0*(1/2)superscript(t/194)
the result was about 62.5,
is this correct.
:
This equation tells you what is left after a period of time, If 80% decayed wouldn't the remainder be .2?
:
I would write it:
A = Ao*.5^(t/h)
Where:
A = .2
Ao = original amt, 1.0 in this case
t = time in days
h = half life in days
:
.2 = 1 * .5^(t/194)
:
ln(.2) = (t/194)*ln(.5)
-1.6094 = (t/194)* -.6931
Which arranges to:
t/194 = -1.6094/-6931
t = 194 * 2.3219
t = 450.45 days for 80% to decay
:
Check it:
A = 1*.5^(450.24/194) = .20001
:
Did this help?